1. Solve the system by addition. 5x – 3y = 13 4x – 3y = 11

2. The sum of two numbers is 90. The second is 10 more than 4 times the first. What are the two numbers?

3. Solve the system by graphing. x + y = 4 –x + y = 2
4. Solve the following system of linear inequalities by graphing
x-y<3
x+2y>6 (these symbols each have lines under them I do not know how to do it)
Solve the system by addition or substitution
3x+6y=0
x=2/3
I still need help with these problems, is there anyone out there that can help me? I will be foreever greatful, thanking you in advance,
nearly crazy

Answer by checkley71(8403)

5X-3Y=13
4X-3Y=11 SUBTRACTING THE SECOND EQUATION FROM THE FIRST WE GET
-----------------
5X-4X=13-11
X=2
5*2-3Y=13
10-3Y=13
-3Y=13-10
-3Y=3
Y=3/-3
Y=-1
PROOF
4*2-3*-1=11
8+3=11
11=11